# Evaluating Zero-Shot Multilingual Aspect-Based Sentiment Analysis with Large Language Models

### v1

@@ Abstract
Aspect-based sentiment analysis (ABSA), a sequence labeling task, has attracted increasing attention in multilingual contexts. While previous research has focused largely on fine-tuning or training models specifically for ABSA, we evaluate large language models (LLMs) under zero-shot conditions to explore their potential to tackle this challenge with minimal task-specific adaptation. We conduct a comprehensive empirical evaluation of a series of LLMs on multilingual ABSA tasks, investigating various prompting strategies, including vanilla zero-shot, chain-of-thought (CoT), self-improvement, self-debate, and self-consistency, across nine different models. Results indicate that while LLMs show promise in handling multilingual ABSA, they generally fall short of fine-tuned, task-specific models. Notably, simpler zero-shot prompts often outperform more complex strategies, especially in high-resource languages like English. These findings underscore the need for further refinement of LLM-based approaches to effectively address ABSA task across diverse languages.The code for experiments and evaluation is available at https://github.com/swaggy66/eval-llm-mabsa.
@@
journal: Knowledge-Based Systems
[…]
[scnu] organization=School of Electronic Science and Engineering (School of Microelectronics),
[…]
@@ 1 Introduction
Aspect-based sentiment analysis (ABSA) is a sequence labeling task in natural language processing (NLP) focused on identifying fine-grained information in sentences, such as aspect terms and their associated sentiment polarity [1]. Although research on ABSA has achieved success on English texts with large corpora, real-world social media interactions often involve multiple languages [2, 3], creating the need for multilingual sentiment analysis.
@@
With the rise of LLMs, there has been growing interest in leveraging in-context learning approaches for various downstream tasks [4, 5, 6, inter alia]. Many studies have evaluated LLMs’ capabilities in sequence labeling tasks such as named entity recognition (NER) [7] and part-of-speech (POS) tagging, particularly in multilingual settings [8, 9]. Similarly, work on LLMs for sentiment analysis has included some assessment of ABSA task [10]. However, few studies have examined LLMs’ performance on multilingual ABSA task specifically through various in-context learning strategies. As a result, the extent to which LLMs can effectively handle multilingual ABSA task under these conditions remains unclear. Understanding how well LLMs capture token-level aspect terms and their associated sentiments across languages will offer insights into their multilingual capabilities and aid in designing more effective models for specific tasks.
@@
2 Background
Multilingual ABSA. Multilingual ABSA aims to identify aspects and their associated sentiments in text across multiple languages. This task is complicated by linguistic diversity, cultural differences, and the scarcity of annotated datasets in non-English languages [11]. Early approaches relied heavily on supervised learning models, necessitating manually annotated data for each language [12, 13, 14, 15]. Recent advancements have shifted focus to cross-lingual embeddings and multilingual pre-trained language models, often fine-tuning these on specific languages for ABSA task [16, 17, 18]. While these approaches have enhanced generalization across languages, they still face difficulties when dealing with user-generated content, such as slang, abbreviations, and culturally dependent expressions [19]. The emergence of LLMs offers new opportunities to tackle these challenges, driving our investigation into their performance on such a nuanced multilingual task.
LLMs for Sequence Labeling tasks. LLMs have recently shown promise in various sequence labeling tasks, such as NER and POS tagging [7, 10, 20]. Unlike traditional models that often require task-specific architectures, LLMs benefit from extensive pre-training on diverse data, enabling them to provide flexible solutions across multiple tasks. Techniques such as prompting and fine-tuning have been explored to improve LLMs’ applicability to these tasks [21, 22, 23]. However, several challenges remain, particularly when adapting LLMs to multilingual contexts and balancing trade-offs between model size, computational cost, and performance [21]. In our work, we extend this research by evaluating LLMs’ effectiveness on multilingual ABSA, a complex task that requires handling both aspect extraction and sentiment classification across languages.
@@ 3 Prompting Strategies
Self-Consistency. The self-consistency prompting strategy aims to stabilize the model’s outputs by running the same experiment multiple times ($n$ iterations) with varying temperature settings. This approach leverages sampling diversity to generate multiple potential outputs, from which the most frequently occurring result is selected as the final answer. The empirical evaluation shows that self-consistency boosts the performance of CoT prompting on a few reasoning benchmarks [22]. This self-consistency is also tested to have improved performance on zero-shot NER tasks [7].
@@ 5.1 Overall Observation
Prompt Sensitivity Across Models, Methods, Languages in Performance. Different prompt designs can lead to significant performance variations. To fairly investigate the impact of this sensitivity on the multilingual ABSA task, we establish uniform performance variations for prompt evaluation for each method. For fine-grained structured outputs, the impact of prompts on performance under the same prompt is significant. Generally, we observe that simple zero-shot prompts lead to better performance than other complex prompts in most cases. In terms of overall performance, GPT-4o achieves the best average performance in all five methods, reaching an F1 score of 45 95%, which can be attributed to the stronger capabilities of the model and its use of a more diverse multilingual corpus during pre-training, leading to better handling of downstream tasks. When comparing the performance between the languages, we clearly observe that the results on English are the best, followed by Spanish and French, and the results on Russian are generally the worst. This is not surprising, as current LLMs are mostly English-centric [9] and we apply English prompts during the experiments. Details about the prompts can be found in the Appendix A.1.
@@ 5.2 Comparison with Fine-Tuning Baseline
5.3 The Effect of Multi-Turn in Performance
As discussed in §5.1, we observe that the performance of both multi-turn approaches, i.e. self-improvement and self-debate is worse than the zero-shot strategies. To better understand this outcome, we analyze the performance across different rounds. Figure 5 presents the average performance differences across all languages in the three rounds for both multi-turn setups. Detailed performance for each language is provided in Appendix A.2.
@@ 5.5 Few-Shot Ablation
In the few-shot setup, the model is provided with $n$ examples from the training data as contextual demonstrations. These examples are selected based on their similarity to the input sentence, as determined using Sentence Transformers [46]. The top $n$ most similar sentences are included as examples, with experiments conducted using the following values for $n$: $n\in[1,2,4,8,16]$.
@@ 5.6 Error Cases
In this section, we analyze the types of errors observed during the experiments. To classify the errors, we follow the categorization proposed by [7], which also addresses token-level tasks like NER:
@@ 6 Conclusion
Appendix A Appendix
@@ A.1 Prompt
We show all of our prompts for six methods. The prompts are in Figure 9, 10 and 11.
@@ A.2 Detailed Results
The details of the overall performance under the self-improvement, self-debate, and few-shot settings are shown in Table 6, 7 and 8.
@@ A.4 Dataset Details
Figure 9: Single-turn prompts used in this paper.
@@
Figure 10: Self-improvement prompts used in this paper.
@@
Figure 11: Self-debate prompts used in this paper.

### v3

@@ Abstract
Aspect-based sentiment analysis (ABSA), a sequence labeling task, has attracted increasing attention in multilingual contexts. While previous research has focused largely on fine-tuning or training models specifically for ABSA, we evaluate large language models (LLMs) under zero-shot conditions to explore their potential to tackle this challenge with minimal task-specific adaptation. We conduct a comprehensive empirical evaluation of a series of LLMs on multilingual ABSA tasks, investigating various prompting strategies, including vanilla zero-shot, chain-of-thought (CoT), self-improvement, self-debate, and self-consistency, across nine different models. Results indicate that while LLMs show promise in handling multilingual ABSA, they generally fall short of fine-tuned, task-specific models. Notably, simpler zero-shot prompts often outperform more complex strategies, especially in high-resource languages like English. These findings underscore the need for further refinement of LLM-based approaches to effectively address ABSA task across diverse languages.
@@ 1 Introduction
Aspect-based sentiment analysis (ABSA) is a sequence labeling task in natural language processing (NLP) focused on identifying fine-grained information in sentences, such as aspect terms and their associated sentiment polarity [1]. Although research [2, 3, 4] on ABSA has achieved success on English texts with large corpora, real-world social media interactions often involve multiple languages [5, 6], creating the need for multilingual sentiment analysis.
@@
With the rise of LLMs, there has been growing interest in leveraging in-context learning approaches for various downstream tasks [7, 8, 9, inter alia]. Many studies have evaluated LLMs’ capabilities in sequence labeling tasks such as named entity recognition (NER) [10] and part-of-speech (POS) tagging, particularly in multilingual settings [11, 12]. Similarly, work on LLMs for sentiment analysis has included some assessment of ABSA task [13]. However, few studies have examined LLMs’ performance on multilingual ABSA task specifically through various in-context learning strategies. As a result, the extent to which LLMs can effectively handle multilingual ABSA task under these conditions remains unclear. Understanding how well LLMs capture token-level aspect terms and their associated sentiments across languages will offer insights into their multilingual capabilities and aid in designing more effective models for specific tasks.
@@
2 Related Work
Multilingual ABSA. Multilingual ABSA aims to identify aspects and their associated sentiment elements in the text across multiple languages. This task is complicated by linguistic diversity, cultural differences, and the scarcity of annotated datasets in non-English languages [14]. Early approaches relied heavily on supervised learning models, necessitating manually annotated data for each language [15, 16, 17, 18]. This reliance on human resources to address data scarcity often made the annotation process extremely costly. Subsequent studies have addressed the limitation of insufficient low-resource sentiment analysis corpora by using machine translation methods. Recent advancements have shifted focus to cross-lingual embeddings and multilingual pre-trained language models, often fine-tuning these on specific languages for ABSA task [19, 20, 21]. Zhang et al. [22] improve the performance of cross-lingual sentiment analysis by introducing code-switched data and knowledge distillation. Lin et al. [23] propose a contrastive learning framework aimed at enhancing the semantic space representation in cross-lingual sentiment analysis tasks. While these approaches have enhanced generalization across languages, they still face difficulties when dealing with user-generated content, such as slang, abbreviations, and culturally dependent expressions [24]. The emergence of LLMs offers new opportunities to tackle these challenges, driving our investigation into their performance on such a nuanced multilingual task.
LLMs for Sequence Labeling tasks. LLMs have recently shown promise in various sequence labeling tasks [25], such as ABSA, NER [26], and POS tagging [10, 13, 27]. Unlike traditional models that often require task-specific architectures, LLMs benefit from extensive pre-training on diverse data, enabling them to provide flexible solutions across multiple tasks. Techniques such as prompting and fine-tuning have been explored to improve LLMs’ applicability to sentiment analysis tasks [28, 29, 30]. Zhong et al. [31] find that the zero-shot performance of LLMs is comparable to that of fine-tuned BERT models. Meanwhile, Wang et al. [32] conduct a preliminary exploration on the application of ChatGPT to open-domain sentiment analysis and reasoning tasks. Han et al. [33] conduct an extensive evaluation of ChatGPT on aspect-based sentiment analysis tasks, revealing a significant performance gap between ChatGPT and other state-of-the-art methods. Deng et al. [34] study the use of pseudo-label data generated by LLMs to fine-tune smaller models, achieving the performance of supervised learning models. However, these efforts have generally been limited to monolingual (English) sentiment analysis tasks, often involving datasets from different domains, which causes the performance of models to fluctuate due to domain-specific differences, complicating the interpretation of these variations. Besides, several challenges remain, particularly when adapting LLMs to multilingual contexts and balancing trade-offs between model size, computational cost, and performance [28]. Exploring LLMs on multilingual parallel corpora can help mitigate data-level differences and allow for a more focused evaluation of model performance across different languages. In our work, we mainly extend this research by evaluating LLMs’ effectiveness on multilingual ABSA, a complex task that requires handling both aspect extraction and sentiment classification across languages.
@@ 3 Prompting Strategies
Self-Consistency. The self-consistency prompting strategy aims to stabilize the model’s outputs by running the same experiment multiple times ($n$ iterations) with varying temperature settings. This approach leverages sampling diversity to generate multiple potential outputs, from which the most frequently occurring result is selected as the final answer, which is called majority voting. The empirical evaluation shows that self-consistency boosts the performance of CoT prompting on a few reasoning benchmarks [29]. This self-consistency is also tested to have improved performance on zero-shot NER tasks [10].
@@ 5.1 Overall Observation
Prompt Sensitivity Across Models, Methods, Languages in Performance. Different prompt designs can lead to significant performance variations. To fairly investigate the impact of this sensitivity on the multilingual ABSA task, we establish uniform performance variations for prompt evaluation for each method. For fine-grained structured outputs, the impact of prompts on performance under the same prompt is significant. Generally, we observe that simple zero-shot prompts lead to better performance than other complex prompts in most cases. In terms of overall performance, GPT-4o achieves the best average performance in all five methods, reaching an F1 score of 45 95%, which can be attributed to the stronger capabilities of the model and its use of a more diverse multilingual corpus during pre-training, leading to better handling of downstream tasks. When comparing the performance between the languages, we clearly observe that the results on English are the best, followed by Spanish and French, and the results on Russian are generally the worst. This is not surprising, as current LLMs are mostly English-centric [12] and we apply English prompts during the experiments. Additionally, the relatively higher performance on Spanish and French can be partially attributed to their linguistic similarity with English, particularly in terms of vocabulary, syntax, and shared Latin roots, which may facilitate transfer learning in multilingual settings. In contrast, Russian is typologically more distant from English, with distinct morphological and syntactic structures, which likely contributes to the lower performance. Moreover, tokenization and subword segmentation may be less optimal for languages with rich morphology like Russian, further degrading model performance. These observations highlight that cross-lingual transfer is not uniform and that language-specific properties play a crucial role in model effectiveness. Details about the prompts can be found in the Appendix A.1.
@@ 5.2 Comparison with Fine-Tuning Baseline
5.3 The Effect of Model Size in Performance
To explore how model size affects performance in ABSA tasks, we conduct experiments using different sizes of the Qwen series models (Qwen2.5 with 0.5B, 1.5B, 7B, 14B, 32B, 72B). In our experimental setup, we focus on representative zero-shot and self-improvement methods in both single-turn and multi-turn dialogues. As shown in Table 3, we observe the following findings: 1) Model performance increases with size. For example, the Qwen2.5 14B model outperforms the Qwen2.5 1.5B model by 21.98% and 22.23% in average F1 score for zero-shot and self-improvement methods, respectively. 2) Zero-Shot performs better overall than Self-Improvement, which aligns with the trends reported in Table 1. In conclusion, larger language models, having been exposed to more data and more complex training frameworks during pretraining, achieve better performance in multilingual ABSA tasks. This also suggests that, for specific downstream tasks, selecting a larger model with superior performance as the base model can be beneficial.
Methods Model EN ES FR NL RU Avg.
Qwen-2.5 0.5B 10.24 8.97 9.88 6.82 4.06 7.99
Qwen-2.5 1.5B 41.26 32.75 25.87 23.01 15.48 27.67
Zero-Shot Qwen-2.5 7B 54.64 48.10 43.60 42.59 37.15 45.22
Qwen-2.5 14B 60.82 52.41 47.77 46.53 40.72 49.65
Qwen-2.5 32B 64.75 55.62 55.46 57.17 42.77 55.50
Qwen-2.5 72B 67.89 59.77 57.36 60.33 46.69 58.41
Qwen-2.5 0.5B 7.88 5.49 6.98 4.68 3.03 5.61
Qwen-2.5 1.5B 33.43 27.65 21.52 17.91 12.33 22.57
Self-Improve. Qwen-2.5 7B 49.14 40.84 38.23 34.72 29.11 38.41
Qwen-2.5 14B 56.47 45.78 42.94 43.13 35.68 44.80
Qwen-2.5 32B 58.39 52.25 50.79 52.08 38.82 50.47
Qwen-2.5 72B 61.35 55.67 54.84 55.37 41.28 53.70
Table 3: The results compared with different sizes of models under zero-shot and self-improvement settings.
5.4 Excursion: Comparison with Triplet Extraction Results
To evaluate the finer-grained aspect-based sentiment analysis triplet extraction task, we additionally extend an assessment of the target-aspect-sentiment detection (TASD) task using Qwen-2.5 7B and GPT-4o, which aims to extract sentiment triplets (aspect, category, sentiment), with the results reported in Table 4. We observe a performance decline across all five prompting strategies after adding the complexity of category recognition (including 12 categories in §A.5). For example, compared to the binary (aspect, sentiment) extraction task, the results on TASD task experience a 3.77% drop in F1 score under the zero-shot strategy. This can be interpreted as the performance of the model deteriorating more noticeably with increasing task complexity. Therefore, it suggests that fine-grained multilingual sentiment triplet extraction tasks require more in-depth exploration and optimization of prompt strategies. Details of the prompts can be found in §A.1.
Methods Model EN ES FR NL RU Avg.
Zero-Shot Qwen-2.5 7B 44.83 41.88 41.66 33.77 33.78 39.18
GPT-4o 48.57 43.28 44.92 39.24 38.75 42.95
CoT Qwen-2.5 7B 42.73 38.28 39.12 31.63 31.38 36.63
GPT-4o 51.34 48.89 46.76 42.14 37.65 45.36
Self-Improve. Qwen-2.5 7B 41.22 36.57 37.49 30.07 29.99 35.07
GPT-4o 46.59 42.37 42.88 37.42 36.84 41.22
Self-Debate Qwen-2.5 7B 37.23 34.57 34.96 28.17 26.16 32.22
GPT-4o 44.23 40.17 41.22 35.84 34.98 39.29
Self-Consist. Qwen-2.5 7B 43.53 40.28 40.31 32.55 32.47 37.83
GPT-4o 49.33 42.44 43.79 38.14 36.95 42.13
Table 4: The triplet extraction results with Qwen-2.5 7B and GPT-4o.
5.5 The Effect of Multi-Turn in Performance
As discussed in §5.1, we observe that the performance of both multi-turn approaches, i.e. self-improvement and self-debate is worse than the zero-shot strategies. To better understand this outcome, we analyze the performance across different rounds. Figure 5 presents the average performance differences across all languages in the three rounds for both multi-turn setups. Overall, the self-improvement setting yields higher average performance compared to the self-debate method, with an increase of 4.82% in the three-round average F1 score. Furthermore, in each individual round, the self-improvement method achieves better performance compared to the self-debate method. Specifically, compared to the first round, the average F1 scores of the self-improvement and self-debate methods decreased by 1.84% and 2.75%, respectively, in the second round. Compared to the second round, the average F1 scores further decreased by 6.67% and 2.74%, respectively, in the third round. Notably, the performance degradation in the self-debate method is more pronounced across multiple dialogue rounds. Detailed performance for each language is provided in Appendix A.2.
@@ 5.7 Few-Shot Ablation
In the few-shot setup, the model is provided with $n$ examples from the training data as contextual demonstrations. A KNN-based retrieval strategy is employed, selecting the top-$n$ nearest examples based on the similarity to the input sentence, as computed using Sentence Transformers [56]. The top $n$ most similar sentences are included as examples, with experiments conducted using the following values for $n$: $n\in[1,2,4,8,16]$.
@@ 5.8 Error Cases
In this section, we analyze the types of errors observed during the experiments. To classify the errors, we follow the categorization proposed by [10], which also addresses token-level tasks like NER. We categorize the prediction errors in sentiment analysis into two main types: aspect term boundaries and sentiment polarity. The error types related to aspect term boundaries include Contain gold, Contained by gold, Overlap with gold, Completely-O, and Omitted mentions. The error types related to sentiment polarity include Out-of-Domain (OOD) types and Wrong types, as shown in the figure below:
@@ 6 Conclusion
[…]
Acknowledgements
The authors would like to respect and thank all reviewers for their constructive and helpful review.
Declarations
  • Funding. This work was supported in part by the Guangdong Basic and Applied Basic Research Foundation under Grant 2023A1515011370, the National Natural Science Foundation of China (32371114), the Characteristic Innovation Projects of Guangdong Colleges and Universities (No. 2018KTSCX049), the Guangdong Provincial Key Laboratory [No.2023B1212060076].
  • Conflict of interest. The authors declare no competing interests.
  • Ethics approval. This atricle does not contain any studies with human participants or animals performed by any of the authors.
  • Data availability. Data sharing is not applicable to this article as no dataset was generated or analyzed during the current study. The dataset used in this paper is public and available.
Appendix A Details of Prompts and Experiments
@@ A.1 Prompt
We show all of our prompts for six methods. The prompts are in Figure 9-14.
@@ A.2 Detailed Results
The details of the overall performance under the self-improvement, self-debate, and few-shot settings are shown in Table 9, 8 and 10.
@@ A.4 Dataset Details
A.5 Detailed Aspect Categories
We present the detailed aspect categories for the dataset in the previous section (§5.4).
Category = [’FOOD#QUALITY’, ’AMBIENCE#GENERAL’, ’LOCATION#GENERAL’, ’SERVICE#GENERAL’, ’RESTAURANT#GENERAL’, ’DRINKS#QUALITY’, ’RESTAURANT#MISCELLANEOUS’, ’DRINKS#PRICES’, ’DRINKS#STYLE_OPTIONS’, ’FOOD#PRICES’, ’FOOD#STYLE_OPTIONS’, ’RESTAURANT#PRICES’]
@@ A.5 Detailed Aspect Categories
Figure 9: Single-turn prompts used in this paper for tuple extraction.
Zero-Shot & Self-Consistency Aspect-Based Sentiment Analysis (ABSA) requires identifying specific entities mentioned in a text and determining the sentiment expressed toward each entity.
Each entity is associated with:
A category from the list: [FOOD#QUALITY, AMBIENCE#GENERAL, LOCATION#GENERAL, SERVICE#GENERAL, RESTAURANT#GENERAL, DRINKS#QUALITY, RESTAURANT#MISCELLANEOUS, DRINKS#PRICES, DRINKS#STYLE_OPTIONS, FOOD#PRICES, FOOD#STYLE_OPTIONS, RESTAURANT#PRICES]
A sentiment: [positive, negative, neutral]. Your task is to: 1. Identify entities in the text, along with their categories and sentiments.
2. For each identified entity, assign a category from the provided category list.
3. Determine the sentiment for each entity as one of [positive, negative, neutral].
4. Return the results as a list of dictionaries, each containing the entity, category, and sentiment. If no entities are found, return an empty list.
Example Output format: [“entity”: “<entity>”, “category”: “<category>”, “sentiment”: “<label>”] Please return the final output based on the following text in json format.
CoT Aspect-Based Sentiment Analysis (ABSA) requires identifying specific entities mentioned in a text and determining the sentiment expressed toward each entity.
Each entity is associated with:
A category from the list: [FOOD#QUALITY, AMBIENCE#GENERAL, LOCATION#GENERAL, SERVICE#GENERAL, RESTAURANT#GENERAL, DRINKS#QUALITY, RESTAURANT#MISCELLANEOUS, DRINKS#PRICES, DRINKS#STYLE_OPTIONS, FOOD#PRICES, FOOD#STYLE_OPTIONS, RESTAURANT#PRICES]
A sentiment: [positive, negative, neutral]. Your task is to: 1. Identify entities in the text, along with their categories and sentiments.
2. For each identified entity, assign a category from the provided category list.
3. Determine the sentiment for each entity as one of [positive, negative, neutral].
4. Provide a reasoning process for how you identified the entities, categories, and sentiments.
5. Return the results as a list of dictionaries, each containing the entity, category, and sentiment. If no entities are found, return an empty list.
Example Output format: [“entity”: “<entity>”, “category”: “<category>”, “sentiment”: “<label>”, “Explanation”: “<reasoning process>”] Please return the final output based on the following text in json format.
Figure 10: Single-turn prompts used in this paper for triplet extraction.
@@
Figure 11: Self-improvement prompts used in this paper for tuple extraction.
Round 1 Aspect-Based Sentiment Analysis (ABSA) requires identifying specific entities mentioned in a text and determining the sentiment expressed toward each entity.
Each entity is associated with:
A category from the list: [FOOD#QUALITY, AMBIENCE#GENERAL, LOCATION#GENERAL, SERVICE#GENERAL, RESTAURANT#GENERAL, DRINKS#QUALITY, RESTAURANT#MISCELLANEOUS, DRINKS#PRICES, DRINKS#STYLE_OPTIONS, FOOD#PRICES, FOOD#STYLE_OPTIONS, RESTAURANT#PRICES]
A sentiment: [positive, negative, neutral]. Your task is to: 1. Identify entities in the text, along with their categories and sentiments.
2. For each identified entity, assign a category from the provided category list.
3. Determine the sentiment for each entity as one of [positive, negative, neutral].
4. Return the results as a list of dictionaries, each containing the entity, category, and sentiment. If no entities are found, return an empty list.
Example Output format: [“entity”: “<entity>”, “category”: “<category>”, “sentiment”: “<label>”] Please return the final output based on the following text in json format.
Round 2 Aspect-Based Sentiment Analysis (ABSA) requires identifying specific entities mentioned in a text and determining the sentiment expressed toward each entity.
Each entity is associated with:
A category from the list: [FOOD#QUALITY, AMBIENCE#GENERAL, LOCATION#GENERAL, SERVICE#GENERAL, RESTAURANT#GENERAL, DRINKS#QUALITY, RESTAURANT#MISCELLANEOUS, DRINKS#PRICES, DRINKS#STYLE_OPTIONS, FOOD#PRICES, FOOD#STYLE_OPTIONS, RESTAURANT#PRICES]
A sentiment: [positive, negative, neutral]. Here is the sentence: {sentence}. You have classified the sentiment and categories of the entities in this sentence. Here is your initial result: {initial_result}. Please explain why you classified them in this way. Example Output format: [“entity”: “<entity>”, “category”: “<category>”, “sentiment”: “<label>”, “Explanation”: “<reasoning process>”] Please return the final output based on the above sentence in json format.
Round 3 Aspect-Based Sentiment Analysis (ABSA) requires identifying specific entities mentioned in a text and determining the sentiment expressed toward each entity.
Each entity is associated with:
A category from the list: [FOOD#QUALITY, AMBIENCE#GENERAL, LOCATION#GENERAL, SERVICE#GENERAL, RESTAURANT#GENERAL, DRINKS#QUALITY, RESTAURANT#MISCELLANEOUS, DRINKS#PRICES, DRINKS#STYLE_OPTIONS, FOOD#PRICES, FOOD#STYLE_OPTIONS, RESTAURANT#PRICES]
A sentiment: [positive, negative, neutral]. Here is the sentence: {sentence}. You have given a sentiment and category classification for the entities in this sentence. Here is your initial result: {initial_result}. Here is your explanation: {explanation_result}. Please recheck your classification and explanation. If you find any errors or better classifications, please update your response accordingly. Example Output format: [“entity”: “<entity>”, “category”: “<category>”, “sentiment”: “<label>”, “Explanation”: “<reasoning process>”] Please return the final output based on the above sentence in json format.
Figure 12: Self-improvement prompts used in this paper for triplet extraction.
@@
Figure 13: Self-debate prompts used in this paper for tuple extraction.
Round 1 Aspect-Based Sentiment Analysis (ABSA) requires identifying specific entities mentioned in a text and determining the sentiment expressed toward each entity.
Each entity is associated with:
A category from the list: [FOOD#QUALITY, AMBIENCE#GENERAL, LOCATION#GENERAL, SERVICE#GENERAL, RESTAURANT#GENERAL, DRINKS#QUALITY, RESTAURANT#MISCELLANEOUS, DRINKS#PRICES, DRINKS#STYLE_OPTIONS, FOOD#PRICES, FOOD#STYLE_OPTIONS, RESTAURANT#PRICES]
A sentiment: [positive, negative, neutral]. Your task is to: 1. Identify the entity with a sentiment mentioned in the given text. 2. For each identified entity, determine the sentiment in the label set (positive, negative, or neutral). 3. The output should be a list of dictionaries, where each dictionary contains the entity with a sentiment and its corresponding sentiment. If there are no sentiment-bearing entities in the text, the output should be an empty list. Example Output format: Round 1: [“entity”: “<entity>”, “category”: “<category>”, “sentiment”: “<label>”] Please return the final output based on the following text in json format.
Round 2 Aspect-Based Sentiment Analysis (ABSA) requires identifying specific entities mentioned in a text and determining the sentiment expressed toward each entity.
Each entity is associated with:
A category from the list: [FOOD#QUALITY, AMBIENCE#GENERAL, LOCATION#GENERAL, SERVICE#GENERAL, RESTAURANT#GENERAL, DRINKS#QUALITY, RESTAURANT#MISCELLANEOUS, DRINKS#PRICES, DRINKS#STYLE_OPTIONS, FOOD#PRICES, FOOD#STYLE_OPTIONS, RESTAURANT#PRICES]
A sentiment: [positive, negative, neutral]. The source sentence is: {sentence}. The first response result: {previous_response} Please review and comment on the following response. Provide corrections if necessary or add more details to improve the result. Example Output format: Round 2: [“entity”: “<entity>”, “category”: “<category>”, “sentiment”: “<label>”,“Review”:“<review>”] Please return the final output based on the above sentence in json format.
Round 3 Aspect-Based Sentiment Analysis (ABSA) requires identifying specific entities mentioned in a text and determining the sentiment expressed toward each entity.
Each entity is associated with:
A category from the list: [FOOD#QUALITY, AMBIENCE#GENERAL, LOCATION#GENERAL, SERVICE#GENERAL, RESTAURANT#GENERAL, DRINKS#QUALITY, RESTAURANT#MISCELLANEOUS, DRINKS#PRICES, DRINKS#STYLE_OPTIONS, FOOD#PRICES, FOOD#STYLE_OPTIONS, RESTAURANT#PRICES]
A sentiment: [positive, negative, neutral]. The source sentence is: {sentence}. The first response result: {first_response}. The first commentary result: {first_commentary}. Based on the initial response and commentary, please further debate and refine the analysis. If there are any conflicting opinions or uncertainties, resolve them (both entity and review) and provide a more detailed and accurate response. Example Output format: Round 3: [“entity”: “<entity>”, “category”: “<category>”, “sentiment”: “<label>”,“Review”:“<review>”] Please return the final output based on the above sentence in json format.
Figure 14: Self-debate prompts used in this paper for triplet extraction.
1 EN 53.37 56.52 60.39 57.10 45.60 48.20 52.19 53.91 39.66 51.88
ES 44.57 48.71 48.80 52.25 32.49 38.34 40.08 46.46 33.79 42.83
FR 43.70 48.74 50.94 49.05 23.15 37.16 39.81 42.94 30.91 40.71
NL 40.76 46.45 50.24 47.73 33.53 33.98 37.41 40.70 26.76 39.73
RU 34.91 40.38 39.34 43.24 30.18 26.59 28.71 34.54 23.79 33.52
Avg. 43.88 48.95 50.94 50.49 34.71 35.91 36.60 42.38 31.95 41.76
2 EN 51.48 52.82 55.83 55.73 41.76 42.86 50.06 52.83 41.15 49.39
ES 43.47 45.59 45.94 50.05 32.07 36.15 40.02 44.72 34.31 41.37
FR 43.10 46.23 49.22 47.40 22.88 33.68 39.83 41.42 32.97 39.64
NL 39.94 41.15 44.73 45.55 32.04 30.03 37.71 37.35 27.50 37.33
RU 34.94 37.44 37.12 40.77 28.80 23.25 28.31 32.10 23.94 31.85
Avg. 42.59 44.65 46.57 47.90 31.51 33.19 39.19 41.69 31.97 39.92
3 EN 51.69 47.69 53.07 50.18 38.83 38.10 48.68 49.14 40.61 46.44
ES 42.69 42.56 43.66 42.83 31.10 31.50 38.29 40.84 34.38 38.65
FR 42.05 42.42 45.70 41.09 21.72 29.80 37.87 38.23 32.75 36.85
NL 38.52 37.53 42.91 37.17 30.71 26.20 34.68 34.72 27.45 34.43
RU 34.26 33.51 35.05 34.69 26.86 21.23 27.03 29.11 23.73 29.50
Avg. 41.84 40.74 44.08 41.19 29.84 29.37 37.31 38.41 31.78 37.17